OpenStudy (jennychan12):

Integral of x/(x^2+1) from 3 to 1?

OpenStudy (jennychan12):

\[\int\limits_{1}^{3}\frac{ x }{ x^2+1 }dx\] ?

OpenStudy (jennychan12):

i set u = x^2 +1 and the answer i got was 1/2 ln 10 but the answer the book showed was 1/2 ln 5 can someone please explain?

OpenStudy (anonymous):

\[\int\limits_{3}^{1}\frac{ x }{ x ^{2}+1 }=\]

OpenStudy (anonymous):

u=x^2+1 so du=2x

OpenStudy (jennychan12):

yeah

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*\int\limits_{3}^{1}\frac{ 1 }{ u }du\]

OpenStudy (anonymous):

that's ln(u)

OpenStudy (anonymous):

1/2*ln(x^2+1) from 3 to 1

OpenStudy (anonymous):

1/2ln(10)-1/2ln(2)

OpenStudy (jennychan12):

1/2 ln u 1/2 ln (x^2+1) from 1 to 3 which would just be 1/2 ln 10 right? the limits are from 1 to 3, not 3 to 1

OpenStudy (anonymous):

so, it is 1/2*ln(10/2)=1/2*ln(5)

OpenStudy (jennychan12):

ohhh wow. i added wrong.

OpenStudy (jennychan12):

thanks anyways :)